# RealGen: Retrieval Augmented Generation for Controllable Traffic Scenarios

### v1

@@ Abstract
Simulation plays a crucial role in the development of autonomous vehicles (AVs) due to the potential risks associated with real-world testing. Although significant progress has been made in the visual aspects of simulators, generating complex behavior among agents remains a formidable challenge. It is not only imperative to ensure realism in the scenarios generated but also essential to incorporate preferences and conditions to facilitate controllable generation for AV training and evaluation. Traditional methods, mainly relying on memorizing the distribution of training datasets, often fall short in generating unseen scenarios. Inspired by the success of retrieval augmented generation in large language models, we present RealGen, a novel retrieval-based in-context learning framework for traffic scenario generation. RealGen synthesizes new scenarios by combining behaviors from multiple retrieved examples in a gradient-free way, which may originate from templates or tagged scenarios. This in-context learning framework endows versatile generative capabilities, including the ability to edit scenarios, compose various behaviors, and produce critical scenarios. Evaluations show that RealGen offers considerable flexibility and controllability, marking a new direction in the field of controllable traffic scenario generation. Check our project website for more information: https://realgen.github.io.
@@ 1 Introduction
Figure 1: (a) Conventional methods make the model memorize the data distribution for generating. (b) In contrast, our method employs a retriever to query datasets (including external data obtained after training) and uses a generative model to generate scenarios by integrating the information from the retrieved scenarios
@@
Retrieval Augmented Generation (RAG) [8, 24], enhancing the generative process by querying related information from external databases, represents great potential in the domain of large language models [36]. In contrast to conventional models that memorize all knowledge within their parameters, as shown in Figure 1(a), RAG models, shown in Figure 1(b), learn to generate comprehensive outputs by retrieving pertinent knowledge from a database, based on the input provided. A notable aspect of RAG is the ability of the database to undergo updates even after the model has been trained, allowing for continual improvement and adaptation. This flexible framework offers possibilities for controllable scenario generation by using appropriate template scenarios as input and facilitating the generation that is not only realistic but also aligned with specific training and evaluation requirements.
In this work, we present RealGen, a retrieval augmented generation framework for generating traffic scenarios. This framework, as shown in Figure 2, initiates with the training of an encoder through contrastive self-supervised learning [41] to enable the retrieval process to query similar scenarios in a latent embedding space. Leveraging this latent representation, we subsequently train a generative model that combines retrieved scenarios to create novel scenarios. The key contributions of this paper are outlined below:
@@ 2.1 Traffic Modeling and Scenario Generation
Research on traffic modeling using generative models has received considerable attention in recent works. Notably, ScenarioNet [37] and Waymax [22] employ large-scale data to train imitation learning models, facilitating the generation of realistic multi-agent scenarios in simulations. SceneGen [48] utilizes the Long Short Term Memory [29] module to autoregressively generate the trajectories for vehicles and pedestrians based on provided maps. Additionally, TrafficSim [47] learns the multi-agent behaviors from real-world data, and TrafficGen [18] proposes a transformer-based autoencoder architecture to accurately model complex interactions involving multiple agents.
Numerous prior studies have employed adversarial generation to synthesize rare yet critical scenarios L2C [15] utilizes a reinforcement learning framework, where the reward for scenario generation is based on the collision rate. To enhance the realism in purely adversarial generation methods, MMG [16] incorporates the data distribution as regularization. Further developments such as AdvSim [53], AdvDO [7] and KING [25] integrate vehicle dynamics to directly optimize the trajectory to find critical scenarios.
@@ 2.2 Retrieval-augmented Generation
One straightforward usage of the retrieved information is to enhance input data through several methods, such as merging the original data and retrieved data [36], employing attention mechanism [3], or extracting skeleton [6]. The rationale behind retrieval systems stems from the impracticality of encoding all knowledge within model parameters, especially considering the dynamic nature of knowledge that evolves with human activities. Consequently, the capability to access external knowledge databases can significantly improve the precision and quality of generated responses, as evidenced in applications involving LLMs [36, 3].
@@ 3 A Latent Representation of Scenario
A crucial element within the retrieval system is the selection metric for data retrieval, which is typically implemented through a distance function between the query sample and candidate samples in the database. Unlike text, which can be converted into word embeddings, traffic scenarios encompass sequential behaviors and intricate interactions among entities, complicating the establishment of a similarity metric for these scenarios. Consequently, in this section, we introduce a scenario autoencoder to extract latent representations that facilitate the assessment of similarity between various traffic scenarios, a vital component in our RAG framework.
@@ 3.1 Scenario Definition
Each scenario is characterized by the trajectories $\tau\in\mathbb{R}^{M\times T\times 5}$, encompassing $M$ agents over a maximum of $T$ time steps. The trajectory of each agent is composed of parameters $[x,y,v,c,s]$, signifying the position $x$, position $y$, velocity $v$, cosine of the heading $c$, and sine of the heading $s$. The initial state of these entities is denoted as $\tau_{0}\in\mathbb{R}^{M\times 5}$. Additionally, the map is encapsulated by $m\in\mathbb{R}^{S\times P\times 4}$, comprised of $S$ lane vectors, with each lane vector consisting of $P$ points. The attributes of these points $[x,y,c,s]$ correspond to the same meaning as those in the agent trajectories.
@@ 3.2 Scenario Autoencoder
Autoencoders [44] are employed to learn compressed latent representations of high-dimensional data, where an encoder projects the data into latent code and a decoder reconstructs the code in the data space. Given that traffic scenarios encompass spatial and temporal dynamics from multiple agents, we design a hierarchical encoder structure alternating between spatial and temporal layers based on transformer architecture [50]. As depicted in the left part of Figure 2, our design incorporates three encoders for behavior, map, and initial position, respectively.
@@
First, we design a behavior encoder ${E}_{b}$ with a spatial-temporal transformer structure that comprises a set of temporal transformer encoders (TransEnc) and a set of spatial transformer encoders. The trajectory $\tau$ is first transformed into a latent embedding $z_{b}$, where $H$ denotes the hidden dimension, through a Multiple Linear Perception (MLP) module. Subsequently, an alternating encoding procedure extracts spatial and temporal features from $z_{b}$. To preserve temporal information, we add a sinusoidal positional embedding (PE) to the embedding prior to employing the temporal transformer. To retain distinct agent information and further compress the behavior feature, we calculate the mean of the latent embedding across the temporal dimension, resulting in the final behavior embedding $z_{b}\in\mathbb{R}^{M\times H}$. Besides the behavior encoder, we also process the map information represented by lane vectors with the attention mechanism [50]. Similarly to the first step in ${E}_{b}$, we project $m$ to the latent space with an MLP module. Then we apply a multi-head attention module (MHA) with layer normalization (LayerNorm) [1] to acquire the map embedding $z_{m}\in\mathbb{R}^{S\times H}$ with a learnable query embedding $q_{m}$.
@@
In developing the decoder, a spatial-temporal transformer architecture was designed to decode the behavior embedding $z_{b}$. Given that $z_{b}$ lacks the temporal dimension, we first replicate it $T$ times to then add a PE before inputting it to the temporal transformer encoder. Throughout the decoding process, the map embedding $z_{m}$ is injected by a cross-attention mechanism:
$z_{b}\leftarrow z_{b}+\text{MHA}(z_{b},z_{m},z_{m}).$ (1)
$\text{MHA}(Q,K,V)$ is the multi-head attention with $Q$, $K$, $V$ representing query, key, and value:
$\begin{split}&\text{MHA}(Q,K,V)=\text{Concatenate}(h_{1},...,h_{i})W^{O}\\ &h_{i}=\text{Attention}(QW^{Q}_{i},KW^{K}_{i},VW^{V}_{i}),\end{split}$ (2)
@@
where $W^{O}$, $W^{Q}_{i}$, $W^{K}_{i}$, and $W^{V}_{i}$ are learnable parameters. The last component of the decoder is an MLP that projects the hidden embedding back into the data domain, resulting in the reconstructed trajectory $\hat{\tau}$. To train the encoders and decoder, mean square error is employed as the reconstruction loss, expressed as $\mathcal{L}_{r}=|\hat{\tau}-\tau|_{2}$. Comprehensive descriptions of the encoding and decoding processes are provided in Algorithm 1.
@@
It is important to note that while the current design of the autoencoder is sufficient for acquiring compressed representations of scenarios, these representations are not invariant to the absolute coordinates and to the order of the agents. This limitation can lead to substantial embedding distances between scenarios that are actually similar. To address these issues, two enhancements are proposed and detailed in the subsequent section.
1 Behavior Encoder $E_{b}$ ($\tau$):
@@
2       for $l_{e}$ in $[1,...,L]$ do
3             $z_{b}\leftarrow$ Temporal TransEnc (PE + $z_{b}$)
4             $z_{b}\leftarrow$ Spatial TransEnc ($z_{b}$)
@@
      $z_{b}\leftarrow$ mean($z_{b}$) // temporal dimension
@@ 9       Initialize a learnable query $q_{m}$
10       $z_{m}\leftarrow\textnormal{LayerNorm (MHA}(q_{m},z_{m},z_{m}))$
11       $z_{m}\leftarrow\textnormal{LayerNorm (}z_{m}+\textnormal{MLP}(z_{m}))$
@@
13
14Initial Pose Encoder $E_{i}$ ($\tau_{0}$):
@@ 18       $z_{r}\leftarrow z_{b}+\textnormal{MHA}(z_{b},z_{i},z_{i})$
19       for $l_{d}$ in $[1,...,L]$ do
20             $z_{r}\leftarrow$ Temporal TransEnc (PE + $z_{r}$)
21             $z_{r}\leftarrow$ Spatial TransEnc ($z_{r}$)
22             $z_{r}\leftarrow z_{r}+\textnormal{MHA}(z_{r},z_{m},z_{m})$
23
@@ 3.3 Invariant feature with contrastive loss
To enhance the representation of scenario similarity in the behavior embedding $z_{b}$, we employ contrastive learning to acquire invariant features. Specifically, we integrate InfoNCE [41] as an additional loss function $\mathcal{L}_{c}$, which optimizes the categorical cross-entropy to distinguish a positive sample from a batch of negative samples. In practice, for a query embedding $z_{b}$, a positive sample $z_{b}^{+}$ is generated by applying random rotation and translation to the original scenario ${\tau,m}$. Meanwhile, negative samples $Z_{b}^{-}$ are selected from the remaining samples in the same batch. In conventional InfoNCE, to bring $z_{b}$ and $z_{b}^{+}$ closer together, the inner product is computed between the embedding $z_{b}$ and the set ${z_{b}^{+},Z_{b}^{-}}$, employing cross-entropy loss to identify the index of the positive sample.
To ensure that the ordering of the agents does not affect the outcomes, $z_{b}\in\mathbb{R}^{M\times H}$ should exhibit permutation invariance across the dimension $M$. Otherwise, merely stacking them into a one-dimensional vector could erroneously represent distinct scenarios as significantly different. To establish permutation invariance, we adopt the Wasserstein distance $W_{2}$ [51] as a more suitable metric than cosine distance for assessing the similarity. By considering $z_{b}$ as a distribution representing $M$ individual behaviors, the intuition behind this choice is that the Wasserstein distance identifies the minimal adjustments necessary for the behaviors in one scenario to mirror those in another. With such a powerful tool, we have the following contrastive loss:
@@
The implementation of this loss results in the behavior embedding $z_{b}$ lacking absolute coordinate data, thereby making the decoder incapable of accurately reconstructing the precise trajectory. To handle this problem, we add the initial poses $\tau_{0}$ of all agents as supplemental input for the decoder. We encode these poses into $z_{i}$ with an MLP encoder $\text{Enc}_{i}$, and then integrate it into the decoder with an MHA module:
$z_{r}\leftarrow z_{b}+\text{MHA}(z_{b},z_{i},z_{i}).$ (4)
Additional specifics regarding the initial pose encoder are detailed in Algorithm 1. In the stage of training, the objective is to minimize a combined loss function, $\mathcal{L}=\mathcal{L}{r}+\lambda\mathcal{L}{c}$, where $\lambda$ serves as a weighting factor for the contrastive loss and is uniformly set to 0.1 across all experiments.
@@ 4.1 The Training of Combiner
Unlike the trajectory prediction task [45] which has ground truth as the optimization target, the objective in training the combiner is to learn the alignment of behaviors in retrieved scenarios with the specified initial pose and map. A well-trained combiner should be able to compose behaviors from these retrieved scenarios, thereby generating new scenarios that have a resemblance to all the retrieved scenarios. This type of objective aligns with the concept of meta-learning, or ‘learning to learn’, as described in [30].
Within our training framework, for a given query scenario ${\tau,m}$ in the dataset, we initially utilize the behavior encoder to obtain the $z_{b}$ and then use KNN to identify $K$ similar behavior embeddings from the database, denoted as $z_{ret}=[z_{b,1},...,z_{b,K}]$. Building on this, we propose a model comprising two MHA modules as the combiner:
@@
where $z_{i}\leftarrow E_{i}(x_{0})$ and $z_{m}\leftarrow E_{m}(m)$ are the initial pose embedding and the map embedding of the query scenario. The gradients are stopped for $z_{i}$ and $z_{m}$. Assuming the $K$ nearest scenarios sufficiently represent the query scenario, it should be feasible to reconstruct the query scenario’s behavior, $z_{b}$, using the retrieved scenario, $z_{rag}$. Consequently, we employ the following loss function as our training goal:
@@
wherein the parameters of the decoder $D\phi$ remain fixed during training. Essentially, this training method is designed to learn an ”inverse” operation of the KNN, aiming to reconstruct a query scenario that closely resembles all $K$ retrieved scenarios in the database.

### v2

@@ Abstract
Simulation plays a crucial role in the development of autonomous vehicles (AVs) due to the potential risks associated with real-world testing. Although significant progress has been made in the visual aspects of simulators, generating complex behavior among agents remains a formidable challenge. It is not only imperative to ensure realism in the scenarios generated but also essential to incorporate preferences and conditions to facilitate controllable generation for AV training and evaluation. Traditional methods, which rely mainly on memorizing the distribution of training datasets, often fail to generate unseen scenarios. Inspired by the success of retrieval augmented generation in large language models, we present RealGen, a novel retrieval-based in-context learning framework for traffic scenario generation. RealGen synthesizes new scenarios by combining behaviors from multiple retrieved examples in a gradient-free way, which may originate from templates or tagged scenarios. This in-context learning framework endows versatile generative capabilities, including the ability to edit scenarios, compose various behaviors, and produce critical scenarios. Evaluations show that RealGen offers considerable flexibility and controllability, marking a new direction in the field of controllable traffic scenario generation. Check our project website for more information: https://realgen.github.io.
@@ 1 Introduction
Figure 1: (a) Conventional methods make the model memorize the data distribution for generating. (b) In contrast, our method employs a retriever to query datasets (including external data obtained after training) and uses a generative model to generate scenarios by integrating the information from the retrieved scenarios.
@@
Retrieval Augmented Generation (RAG) [8, 25], which enhances the generative process by querying related information from external databases, represents great potential in the domain of large language models [37]. In contrast to conventional models that memorize all knowledge within their parameters, as shown in Figure 1(a), RAG models, shown in Figure 1(b), learn to generate comprehensive outputs by retrieving pertinent knowledge from a database, based on the input provided. A notable aspect of RAG is the ability of the database to undergo updates even after the model has been trained, allowing continuous improvement and adaptation. This flexible framework offers possibilities for controllable scenario generation by using appropriate template scenarios as input and facilitating the generation that is not only realistic but also aligned with specific training and evaluation requirements.
In this work, we present RealGen, a retrieval augmented generation framework for generating traffic scenarios. This framework, as shown in Figure 2, begins with the training of an encoder through contrastive self-supervised learning [42] to allow the retrieval process to query similar scenarios in a latent embedding space. Leveraging this latent representation, we subsequently train a generative model that combines retrieved scenarios to create novel scenarios. The key contributions of this paper are summarized below.
@@ 2.1 Traffic Modeling and Scenario Generation
Research on traffic modeling using generative models has received considerable attention in recent works. Notably, ScenarioNet [38] and Waymax [23] employ large-scale data to train imitation learning models, facilitating the generation of realistic multi-agent scenarios in simulations. SceneGen [50] utilizes the Long Short Term Memory [30] module to autoregressively generate the trajectories for vehicles and pedestrians based on provided maps. Additionally, TrafficSim [47] learns the multi-agent behaviors from real-world data, and TrafficGen [18] proposes a transformer-based autoencoder architecture to accurately model complex interactions involving multiple agents. MixSim [48] builds a reactive digital twin and finds safety-critical scenarios with black-box optimization. RTR [61] models reactive behaviors of vehicles using the combination of reinforcement learning and imitation learning.
Numerous prior studies have employed adversarial generation to synthesize rare yet critical scenarios. L2C [17] utilizes a reinforcement learning framework, where the reward for scenario generation is based on the collision rate. To enhance the realism in purely adversarial generation methods, MMG [15] incorporates the data distribution as regularization. Further developments such as AdvSim [54], AdvDO [7], and KING [26] integrate vehicle dynamics to directly optimize the trajectory to find critical scenarios.
@@ 2.2 Retrieval-augmented Generation
One intuitive usage of the retrieved information is to enhance input data through several methods, such as merging the original data and retrieved data [37], employing attention mechanism [3], or extracting skeleton [6]. The rationale behind retrieval systems stems from the impracticality of encoding all knowledge within model parameters, especially considering the dynamic nature of knowledge that evolves with human activities. Consequently, the capability to access external knowledge databases can significantly improve the precision and quality of generated responses, as evidenced in applications involving LLMs [37, 3].
@@ 3 A Latent Representation of Scenario
A crucial element within the retrieval system is the selection metric for data retrieval, which is typically implemented through a distance function between the query sample and the candidate samples in the database. Unlike text, which can be converted into word embeddings, traffic scenarios encompass sequential behaviors and intricate interactions among entities, complicating the establishment of a similarity metric for these scenarios. Consequently, in this section, we introduce a scenario autoencoder to extract latent representations that facilitate the assessment of similarity between various traffic scenarios, a vital component in our RAG framework.
@@ 3.1 Scenario Definition
Each scenario is characterized by the trajectories $\tau\in\mathbb{R}^{M\times T\times 5}$, encompassing $M$ agents over a maximum of $T$ time steps. The trajectory of each agent is composed of parameters $[x,y,v,c,s]$, signifying position $x$, position $y$, velocity $v$, cosine of the heading $c$, and sine of the heading $s$. The initial state of these entities is denoted as $\tau_{0}\in\mathbb{R}^{M\times 5}$. Furthermore, the map is encapsulated by $m\in\mathbb{R}^{S\times 4}$, composed of $S$ lane segments. The attributes of these points $[x_{s},y_{s},x_{e},y_{e}]$ correspond to the starting and ending positions of each segment.
@@ 3.2 Scenario Autoencoder
Autoencoders [44] are used to learn compressed latent representations of high-dimensional data, where an encoder projects the data into latent code and a decoder reconstructs the code in the data space. Given that traffic scenarios encompass spatial and temporal dynamics from multiple agents, we design a hierarchical encoder structure alternating between spatial and temporal layers based on transformer architecture [51]. As depicted on the left of Figure 2, our design incorporates three encoders for behavior, map, and initial position.
@@
First, we design a behavior encoder ${E}_{b}$ with a spatial-temporal transformer structure that comprises $L_{e}$ temporal transformer encoders ($\text{Encoder}_{t}^{i}$) and $L_{e}$ spatial transformer encoders ($\text{Encoder}_{s}^{i}$), where $i\in\{1,\dots,L_{e}\}$. The trajectory $\tau$ is first transformed into a latent embedding $z_{b}$, where $H$ denotes the hidden dimension, through a Multiple Linear Perception (MLP) module. Subsequently, an alternating encoding procedure extracts spatial and temporal features from $z_{b}$. To preserve temporal information, we add a sinusoidal positional embedding (PE) to the embedding before employing the temporal transformer. To retain distinct agent information and further compress the behavior feature, we calculate the mean of the latent embedding across the temporal dimension, resulting in the final behavior embedding $z_{b}\in\mathbb{R}^{M\times H}$. Besides the behavior encoder, we also process the map information represented by lane vectors with the attention mechanism [51]. Similarly to the first step in ${E}_{b}$, we project $m$ to the latent space with an MLP module. Then we apply a multi-head attention module (MHA) with layer normalization (LN) [1] to acquire the map embedding $z_{m}\in\mathbb{R}^{S\times H}$ with a learnable query embedding $q_{m}$.
@@
In the decoder, a spatial-temporal transformer architecture, similar to the behavior encoder, is designed to obtain the behavior embedding $z_{b}$ through $L_{d}$ times encoding. Given that $z_{b}$ lacks the temporal dimension, we first replicate it $T$ times and then add a PE before entering it into the temporal transformer encoder. Throughout the decoding process, the map embedding $z_{m}$ is injected by a cross-attention mechanism: $z_{b}\leftarrow z_{b}+\text{MHA}(z_{b},z_{m},z_{m}).$ $\text{MHA}(Q,K,V)$ is the multi-head attention with $Q$, $K$, $V$ representing query, key, and value:
$\begin{split}&\text{MHA}(Q,K,V)=\text{Concatenate}(h_{1},...,h_{i})W^{O},\\ &h_{i}=\text{Attention}(QW^{Q}_{i},KW^{K}_{i},VW^{V}_{i}),\end{split}$ (1)
@@
where $W^{O}$, $W^{Q}_{i}$, $W^{K}_{i}$, and $W^{V}_{i}$ are learnable parameters. The last component of the decoder is an MLP that projects the hidden embedding back into the data domain, resulting in the reconstructed trajectory $\hat{\tau}$. To train the encoders and decoder, the mean square error is used as the reconstruction loss, expressed as $\mathcal{L}_{r}=|\hat{\tau}-\tau|_{2}$. Comprehensive descriptions of the encoding and decoding processes are provided in Algorithm 1.
@@
It is important to note that, while the current design of the autoencoder is sufficient for acquiring compressed representations of scenarios, these representations are not invariant to the absolute coordinates and the order of the agents. This limitation can lead to substantial embedding distances between similar scenarios. To address these issues, two enhancements are proposed and detailed in the following section.
1
Behavior Encoder $E_{b}$ ($\tau$):
@@
2       for $i$ in $[1,...,L_{e}]$ do
3             $z_{b}\leftarrow$ $\text{Encoder}_{t}^{i}$ (PE + $z_{b}$)
4             $z_{b}\leftarrow$ $\text{Encoder}_{s}^{i}$ ($z_{b}$)
@@
6      $z_{b}\leftarrow$ mean ($z_{b}$)
@@ 10       Initialize a learnable query $q_{m}$
11       $z_{m}\leftarrow\textnormal{LN (MHA}\ (q_{m},z_{m},z_{m}))$
12       $z_{m}\leftarrow\textnormal{LN (}z_{m}+\textnormal{MLP}\ (z_{m}))$
@@
14
15Initial Pose Encoder $E_{i}$ ($\tau_{0}$):
@@ 19       $z_{r}\leftarrow z_{b}+\textnormal{MHA}(z_{b},z_{i},z_{i})$
20       for $i$ in $[1,...,L_{d}]$ do
21             $z_{r}\leftarrow$ $\text{Encoder}_{t}^{i}$ (PE + $z_{r}$)
22             $z_{r}\leftarrow$ $\text{Encoder}_{s}^{i}$ ($z_{r}$)
23             $z_{r}\leftarrow z_{r}+\textnormal{MHA}^{i}(z_{r},z_{m},z_{m})$
24
@@ 3.3 Invariant feature with contrastive loss
To enhance the representation of scenario similarity in the behavior embedding $z_{b}$, we employ contrastive learning to acquire invariant features. Specifically, we integrate InfoNCE [42] as an additional loss function $\mathcal{L}_{c}$, which optimizes categorical cross-entropy to distinguish a positive sample from a batch of negative samples. In practice, for a query embedding $z_{b}$, a positive sample $z_{b}^{+}$ is generated by applying random rotation and translation to the original scenario ${\tau,m}$. Meanwhile, negative samples $Z_{b}^{-}$ are selected from the remaining samples in the same batch. In conventional InfoNCE, to bring $z_{b}$ and $z_{b}^{+}$ closer together, the inner product is calculated between the embedding $z_{b}$ and the set ${z_{b}^{+},Z_{b}^{-}}$, employing cross-entropy loss to identify the index of the positive sample.
To ensure that the ordering of the agents does not affect the outcomes, $z_{b}\in\mathbb{R}^{M\times H}$ should exhibit permutation invariance across the dimension $M$. Otherwise, simply stacking them into a one-dimensional vector could erroneously represent distinct scenarios as significantly different. To establish permutation invariance, we adopt the Wasserstein distance $W_{2}$ [52] as a more suitable metric than cosine distance for assessing the similarity. Taking $z_{b}$ as a distribution representing $M$ individual behaviors, the intuition behind this choice is that the Wasserstein distance identifies the minimal adjustments necessary for behaviors in one scenario to reflect those in another. With such a powerful tool, we have the following contrastive loss:
@@
The implementation of this loss results in the behavior embedding $z_{b}$ lacking absolute coordinate data, making the decoder incapable of accurately reconstructing the precise trajectory. To handle this problem, we add the initial poses $\tau_{0}$ of all agents as a supplementary input for the decoder. We encode these poses into $z_{i}$ with an MLP encoder $\text{Enc}_{i}$, and then integrate it into the decoder with an MHA module $z_{r}\leftarrow z_{b}+\text{MHA}(z_{b},z_{i},z_{i})$. Additional specifics regarding the initial pose encoder are detailed in Algorithm 1. In the training stage, the objective is to minimize a combined loss function $\mathcal{L}=\mathcal{L}_{r}+\lambda\mathcal{L}_{c}$, where $\lambda$ serves as a weighting factor for contrastive loss and is uniformly set at 0.1 for all experiments.
@@ 4.1 The Training of Combiner
Unlike the trajectory prediction task [45] which has ground truth as the optimization target, the objective in training the combiner is to learn the alignment of behaviors in the retrieved scenarios with the initial pose and the specified map. A well-trained combiner should be able to compose behaviors from these retrieved scenarios, thereby generating new scenarios that have a resemblance to all the retrieved scenarios. This type of objective aligns with the concept of meta-learning, or ‘learning to learn’, as described in [31].
Within our training framework, for a given query scenario ${\tau,m}$ in the dataset, we initially utilize the behavior encoder to obtain $z_{b}$ and then use KNN to identify $K$ similar behavior embeddings from the database, denoted $z_{ret}=[z_{b,1},...,z_{b,K}]$. Building on this, we propose a model comprising two MHA modules as the combiner:
@@
where $z_{i}\leftarrow E_{i}(x_{0})$ and $z_{m}\leftarrow E_{m}(m)$ are the initial pose embedding and the map embedding of the query scenario. Gradients are stopped for $z_{i}$ and $z_{m}$. Assuming that the $K$ nearest scenarios adequately represent the query scenario, it should be feasible to reconstruct the behavior of the query scenario, $z_{b}$, using the retrieved scenario, $z_{rag}$. Consequently, we employ the following loss function as our training goal:
@@
wherein the parameters of the decoder $D\phi$ remain fixed during training. Essentially, this training method is designed to learn an "inverse" operation of the KNN, aiming to reconstruct a query scenario that closely resembles all $K$ retrieved scenarios in the database.